OpenStudy (anonymous):

Can anyone please help me with the following? I can't remember from my high school algebra... Square root of (1+9x^2) I can't remember how to simplify the term to be able to square root it easily, because I know you may not say 1+9x^2 = (1+3x)^2

ganeshie8 (ganeshie8):

\(\large \sqrt{1+9x^2}\)

ganeshie8 (ganeshie8):

trying to see, if i understood the question correctly :)

ganeshie8 (ganeshie8):

we cant simplify it right any further, right ?

ganeshie8 (ganeshie8):

btw, below statement doesn't look corrct :- \(1+9x^2 = (1+3x)^2 \)

OpenStudy (anonymous):

Yes, that's exactly what it looks like :)

OpenStudy (anonymous):

And yes, I assumed I may not do that assumption :)

ganeshie8 (ganeshie8):

okay :) lets see if below is correct or not :- \(1+9x^2=(1+3x)^2\)

ganeshie8 (ganeshie8):

how can we **know** if above equation we have put is correct or not ? shall we expand the right side and see ? :)

ganeshie8 (ganeshie8):

do u remember this formula : \((a+b)^2 = a^2+b^2+2ab\)

OpenStudy (anonymous):

Yes, then we get \[1 + 6x + 9x^2?\]

ganeshie8 (ganeshie8):

yup :) is it fine to conclude \(1+9x^2 \ne (1+3x)^2\) ?

OpenStudy (anonymous):

Definitely :) Is there then no way to simplify \[1 + 9x^2\]? :(

ganeshie8 (ganeshie8):

\(\sqrt{1+9x^2}\) i don't think we can do anything much here. it looks fully simplified to me.... :|

OpenStudy (anonymous):

Awww, dangit. Thank so much anyways :)

ganeshie8 (ganeshie8):

may i knw the context of this q, so we can discuss further... if u want ? :)

OpenStudy (anonymous):

Sure! It was the following: Find the surface area generated when the graph of the equation \[f(x)=x^3\] is revolved around the x-axis at points A(1,1) and B(2,8)

ganeshie8 (ganeshie8):

Ahh calculus !

ganeshie8 (ganeshie8):

then im sure u knw pretty well wat u doing lol ^_^

OpenStudy (anonymous):

And when I differentiated \[x^3\] I got \[3x^2\] and in my formula for getting the surface area, i had to solve \[\sqrt{1 +(3x^2)^2}\]Maybe I missed something???

OpenStudy (anonymous):

Hahaha thanks ^_^

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):